# Multiscale Training of Convolutional Neural Networks

### v1

@@ Appendix D Additional Results
D.1 Experiments for the denoising task on the CelebA dataset
To observe the behavior of the Full Multiscale algorithm, we performed additional experiments for the denoising task on the CelebA dataset using UNet, ResNet, MFC-UNet, and MFC-ResNet. The results have been presented in Table 6 showing that both multiscale training strategies achieved similar performance to SGD for all networks with a considerably lower number of #WU .
@@ D.1 Experiments for the denoising task on the CelebA dataset
Table 6: Comparison between the performance of different networks (UNet, ResNet, MFC-UNet, and MFC-ResNet) for the denoising task on the CelebA dataset using various training strategies. We report the mean MSE computed over 512 images from the validation set and the total #WU under each training framework.
D.2 Experiments and comparisons with deeper networks
@@
Table 7: Comparison of training strategies on the denoising task on the STL10 dataset for ResNet18, ResNet50 and a deeper UNets.
D.3 Comparison with Fourier Base Convolutions in Super-Resolution Tasks
We conducted experiments on the Urban100 dataset to evaluate the performance of different convolutional techniques in super-resolution tasks. Specifically, we compared the widely used ESPCN architecture with our proposed MFC-ESPCN (which uses parabolic, mesh-free convolutions) and FNO-ESPCN (which incorporates spectral convolutions used by the FNO paper). The results, measured using the SSIM metric, highlight that our MFC-ESPCN achieves comparable or better performance than standard convolutional methods while requiring lower computational resources. Furthermore, MFC-ESPCN outperforms the FNO-based spectral convolution approach, as it effectively captures both local and global dependencies without being limited by the truncation of frequencies in the Fourier domain. The results for the task of multi-resolution are presented in Table 8
Training Strategy SSIM ($\uparrow$)
ESPCN MFC-ESPCN FNO-ESPCN
SGD (Single Scale) 0.841 0.842 0.744
Multiscale-SGD (Ours) 0.822 0.831 0.743
Full Multiscale-SGD (Ours) 0.811 0.823 0.732
Table 8: Performance Comparison of ESPCN, MFC-ESPCN, and FNO-ESPCN across different training strategies, on the super-resolution task on the Urban100 dataset. SSIM is used as the evaluation metric, where higher values indicate better image similarity to the ground truth.
D.4 Experiments with a fixed budget
An additional way to observe convergence is to plot the convergence as a function of work units. In this way, it is possible to assess the convergence of different methods for giving some value to the work units. To this end, we plot such a curve in Figure 5.
Figure 5: Convergence of different methods as a function of work-units
D.5 Training for longer
We have experimented with longer training of our network by doubling the number of epochs of the experiment reported in Table 7 for the ResNet-18 and ResNet-50. The results for a single mesh (SGD) are summarized in Table 9.
Double #iterations Architecture MSE ($\downarrow$)
No ResNet-18 0.1622
No ResNet-50 0.1611
Yes ResNet-18 0.1599
Yes ResNet-50 0.1585
Table 9: Experiments with longer training (double the number of iterations) of ResNet18 and ResNet50 on the denoising task on the STL10 dataset.
D.6 Comparison of runtimes between a single standard 2D convolution and MFC
@@
Image Size (n×$\times$×n) / Kernel Size (k×$\times$×k) 32 64 128 256
5 $2.62\times 10^{7}/5.56\times 10^{7}$ $1.05\times 10^{8}/2.65\times 10^{8}$ $4.19\times 10^{8}/1.23\times 10^{9}$ $1.68\times 10^{9}/5.57\times 10^{9}$
7 $5.14\times 10^{7}/5.56\times 10^{7}$ $2.06\times 10^{8}/2.65\times 10^{8}$ $8.22\times 10^{8}/1.23\times 10^{9}$ $3.29\times 10^{9}/5.57\times 10^{9}$
9 $8.49\times 10^{7}/5.56\times 10^{7}$ $3.40\times 10^{8}/2.65\times 10^{8}$ $1.36\times 10^{9}/1.23\times 10^{9}$ $5.44\times 10^{9}/5.57\times 10^{9}$
11 $1.27\times 10^{8}/5.56\times 10^{7}$ $5.08\times 10^{8}/2.65\times 10^{8}$ $2.03\times 10^{9}/1.23\times 10^{9}$ $8.12\times 10^{9}/5.57\times 10^{9}$
13 $1.77\times 10^{8}/5.56\times 10^{7}$ $7.09\times 10^{8}/2.65\times 10^{8}$ $2.84\times 10^{9}/1.23\times 10^{9}$ $1.13\times 10^{10}/5.57\times 10^{9}$
15 $2.36\times 10^{8}/5.56\times 10^{7}$ $9.44\times 10^{8}/2.65\times 10^{8}$ $3.77\times 10^{9}/1.23\times 10^{9}$ $1.51\times 10^{10}/5.57\times 10^{9}$
Table 10: Comparison of FLOPs for Standard Convolutions and MFCs across different image and kernel sizes. The reported FLOPs are in the format of Standard Convolutions / our MFCs.
D.7 Multiscale-SGD vs. Random Crops
A method that implicitly includes multiscale information during training is that of the data-augmentation technique of random crops. We now compare the performance between using random crops and our Multiscale-SGD, using the ’RandomResizedCrop’ transform from torchvision in PyTorch (Paszke et al., 2017). As shown in Table 11, the performance with random crops exhibits a slight deterioration compared to fixed-size images. This reduction in performance can likely be attributed to the smaller receptive field introduced by the use of random crops, compared with our Multiscale-SGD, which uses multiple resolutions of the image.
Training Strategy UNet MFC (Random crops) UNet Regular Convs. (Random crops) UNet MFC UNet Regular Convs.
Multiscale 0.1623 0.1612 0.1605 0.1604
Full Multiscale 0.1619 0.1601 0.1609 0.1597
Single Mesh 0.1602 0.1621 0.1594 0.1610
Table 11: Comparison of training strategies with UNet using MFCs and regular convolutions, with random crops vs. our Multiscale-SGD.
@@ Appendix E Experimental Setting
For the denoising task, the experiments were conducted on STL10 and CelebA datasets. The networks used were UNet, ResNet with standard convolutions MFC-UNet, and MFC-ResNet with mesh-free convolutions. The key details of the experimental setup are summarized in Table 12. For the image super-resolution task, the experiments were conducted on the Urban100 dataset, employing a deep residual network and utilizing a multiscale training strategy. The key details of the experimental setup, including the optimizer, loss function, and training strategies, are summarized in Table 13. All our experiments were conducted on an NVIDIA A6000 GPU with 48GB of memory. Upon acceptance, we will release our source code, implemented in PyTorch (Paszke et al., 2017).
Table 12: Experimental details for training for the denoising task
@@
UNet: 2,537,187;
ResNet: 597,699,
MFC-UNet: 2,307,411,
@@
SGD and Multiscale-SGD: [2000, 2000, 2000, 2000],
Full-Multiscale-SGD: [2000, 1000, 500, 250] for the 4 levels
Evaluation metrics MSE over the 512 images from the validation set
Table 13: Experimental details for training for the super-resolution task
@@
Dataset Urban100, consisting of paired low-resolution and high-resolution image patches extracted for training and validation.
Network Architecture A deep residual network (ResNet) with 5 residual blocks, specifically designed for image super-resolution. The ESPCN and the modified architectures with MFC were also included for a better comparison.
Training Strategy A multiscale training approach is employed, utilizing Full-Multiscale-SGD cycles to progressively refine image resolution during training.
Loss Function The Structural Similarity Index Measure (SSIM) is used to maximize perceptual quality between predicted and target images.
Optimizer The Adam (Kingma & Ba, 2014) optimizer, initialized with a learning rate of 1e-3, is used for its adaptive gradient handling.
Learning Rate Set to an initial value of 1e-3.
Batch Size Strategy Dynamic batch sizing is used, adjusting the batch size upwards during different stages of training for improved efficiency. For details, see Appendix C.
Multiscale Levels A hyperparameter capped at 4.
Early Stopping Implemented based on the validation loss and gradient norms, aiming to prevent overfitting and reduce computation time.
Evaluation Metric Validation loss is calculated using SSIM to evaluate the quality of the generated high-resolution images.
Appendix F Runtimes
To have a fair time comparison, we measure the time for the network to process the same number of examples in each of the experiments. Let us consider the case of $4$ levels within a Multiscale-SGD training strategy. In this case, a multiscale computation of the gradient processes for every single image on the finest resolution, 2 images on the next resolution, $4$ on the third resolution, and $8$ on the coarsest resolution. Thus, we compare its runtimes with the time it takes to process a batch of $15$ fine-scale images, as with the standard approach of SGD. The results are measured on ResNet and MFC-ResNet architectures.
We observe a significant time reduction when considering the multiscale strategy. Note that the MFC-ResNet is more expensive than the standard ResNet. This is due to the computation of the Fourier transform, which is more expensive than a simple convolution.
ResNet
Training Strategy SGD Multiscale-SGD (Ours, 4 levels)
Time (ms) 21.5 1.43
MFC-ResNet
Training Strategy SGD Multiscale-SGD (Ours, 4 levels)
Time (ms) 27.3 1.79
Table 14: Runtime (milliseconds) per training iteration using standard convolution kernels, and our MFCs. The results are measured on ResNet and MFC-ResNet architectures with an input image size of $256\times 256$ on an NVIDIA A6000 GPU.
Appendix G Visualizations
@@
In this section, we include visualization of the outputs obtained from baseline UNet and ResNet models, as well as our MFC-UNet and MFC-ResNet networks. The visualizations for the super-resolution task are provided in Figure 11 and Figure 12. The visualizations for the denoising task on the STL-10 dataset are provided in Figure 7 and Figure 8. The visualizations for the denoising task on the CelebA dataset are provided in Figure 9 and Figure 10.
Figure 7: A comparison of different network predictions for Single Scale, Multiscale, and Full-Multiscale for an image from the STL10 dataset. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 8: A comparison of different network predictions for Single Scale, Multiscale, and Full-Multiscale for an image from the STL10 dataset. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 9: A comparison of different network predictions for Single Scale, Multiscale, and Full-Multiscale for an image from the CelebA dataset. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 10: A comparison of different network predictions for Single Scale, Multiscale, and Full-Multiscale for an image from the CelebA dataset. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 11: A comparison of different networks predictions for SGD, Multiscale-SGD, and Full- Multiscale-SGD for an image from the Urban100 dataset.
Figure 12: A comparison of different networks predictions for Single Scale, Multiscale, and Full- Multiscale for an image from the Urban100 dataset..
Appendix H Comparing feature maps between CNNs and MFCs
@@
Appendix I Analysis of the cost of training compared to loss
One possible way to observe the advantage of multiscale training is to record the loss as a function of computational budget; that is, rather than recording the number of iterations, we record the loss as a function of computational cost. To this end we add a figure of the loss as a function of WU.

### v2

@@ Appendix D Additional results
D.1 Experimental details and training convergence analysis for the blur estimation task
(a) Standard Convolution
(b) Mesh-Free Convolution
(c) Training Loss
For the blur estimation task, we assume that both the image ${\bf u}$ and its blurred version ${\bf y}$ are available on a fine mesh of size $256^{2}$. The relationship is linear, given by ${\bf y}={\bf G}\star{\bf u}+{\bf z}$, where ${\bf G}$ is the blurring kernel. Given ${\bf u}$ and ${\bf y}$, we solve the optimization problem $\min_{{\bf G}}{\frac{1}{2}}{\mathbb{E}}_{{\bf u},{\bf y}}\|{\bf y}-{\bf G}\star{\bf u}\|^{2}.$
This is a quadratic stochastic optimization problem for the kernel ${\bf G}$. We simulate ${\bf u}$ as a Random Markov Field (Tenorio et al., 2011) using a random $7\times 7$ blurring kernel for ${\bf G}$. The problem is solved with standard CNNs and MFCs, comparing SGD, Multiscale-SGD, and Full-Multiscale-SGD strategies. As seen in Figure 4, for standard convolutions, the loss increases with mesh refinement, making multiscale approaches ineffective. However, our MFCs maintain stable loss across meshes, making the multiscale process efficient, with the final 50 iterations on the finest mesh.
D.2 Experiments for the denoising task on the CelebA dataset
To observe the behavior of the Full Multiscale algorithm, we performed additional experiments for the denoising task on the CelebA dataset using UNet, ResNet, MFC-UNet, and MFC-ResNet. The results have been presented in Table 5, showing that both multiscale training strategies achieved similar performance to SGD for all networks with a considerably lower number of #WU .
Table 5: Comparison of different training strategies using UNet, ResNet, MFC-UNet, and MFC-ResNet for the denoising task on the CelebA dataset.
@@ D.2 Experiments for the denoising task on the CelebA dataset
D.3 Experiments with deeper networks
Table 6: Comparison of different training strategies using standard convolution and MFC-based variants of ResNet18, ResNet50, and a deeper UNet for the denoising task on the STL10 dataset.
@@ D.3 Experiments with deeper networks
D.4 Comparison with Fourier Base Convolutions for the super-resolution task
Table 7: Comparison of different training strategies using ESPCN, MFC-ESPCN, and FNO-ESPCN for the super-resolution task on the Urban100 dataset.
Training strategy SSIM ($\uparrow$) ESPCN MFC-ESPCN FNO-ESPCN SGD (Single Scale) 0.841 0.842 0.744 Multiscale-SGD (Ours) 0.822 0.831 0.743 Full-Multiscale-SGD (Ours) 0.811 0.823 0.732
We conducted experiments on the Urban100 dataset to evaluate the performance of different convolutional techniques for the super-resolution task. Specifically, we compared the widely used ESPCN architecture with our proposed MFC-ESPCN (which uses parabolic, mesh-free convolutions) and FNO-ESPCN (which incorporates spectral convolutions used by the FNO paper). The results, measured using the SSIM metric, highlight that our MFC-ESPCN achieves comparable or better performance than standard convolutional methods while requiring lower computational resources. Furthermore, MFC-ESPCN outperforms the FNO-based spectral convolution approach, as it effectively captures both local and global dependencies without being limited by the truncation of frequencies in the Fourier domain. The results for the task of multi-resolution are presented in Table 7.
D.5 Comparison of runtimes between a standard convolution and MFC and between SGD and Multiscale-SGD
@@
Table 8: Comparison of FLOPs for standard convolutions and MFCs across different image and kernel sizes. The reported FLOPs are in the format of FLOPs for standard convolutions / FLOPs for our MFCs.
Kernel size $(k\times k)\rightarrow$
Image size $(n\times n)\downarrow$
$k=32$ $k=64$ $k=128$ $k=256$
$n=5$ 2.62 ×$\times 10^{7}$ / 5.56 ×$\times 10^{7}$ 1.05 ×$\times 10^{8}$ / 2.65 ×$\times 10^{8}$ 4.19 ×$\times 10^{8}$ / 1.23 ×$\times 10^{9}$ 1.68 ×$\times 10^{9}$ / 5.57 ×$\times 10^{9}$
$n=7$ 5.14 ×$\times 10^{7}$ / 5.56 ×$\times 10^{7}$ 2.06 ×$\times 10^{8}$ / 2.65 ×$\times 10^{8}$ 8.22 ×$\times 10^{8}$ / 1.23 ×$\times 10^{9}$ 3.29 ×$\times 10^{9}$ / 5.57 ×$\times 10^{9}$
$n=9$ 8.49 ×$\times 10^{7}$ / 5.56 ×$\times 10^{7}$ 3.40 ×$\times 10^{8}$ / 2.65 ×$\times 10^{8}$ 1.36 ×$\times 10^{9}$ / 1.23 ×$\times 10^{9}$ 5.44 ×$\times 10^{9}$ / 5.57 ×$\times 10^{9}$
$n=11$ 1.27 ×$\times 10^{8}$ / 5.56 ×$\times 10^{7}$ 5.08 ×$\times 10^{8}$ / 2.65 ×$\times 10^{8}$ 2.03 ×$\times 10^{9}$ / 1.23 ×$\times 10^{9}$ 8.12 ×$\times 10^{9}$ / 5.57 ×$\times 10^{9}$
$n=13$ 1.77 ×$\times 10^{8}$ / 5.56 ×$\times 10^{7}$ 7.09 ×$\times 10^{8}$ / 2.65 ×$\times 10^{8}$ 2.84 ×$\times 10^{9}$ / 1.23 ×$\times 10^{9}$ 1.13 ×$\times 10^{1}0$ / 5.57 ×$\times 10^{9}$
$n=15$ 2.36 ×$\times 10^{8}$ / 5.56 ×$\times 10^{7}$ 9.44 ×$\times 10^{8}$ / 2.65 ×$\times 10^{8}$ 3.77 ×$\times 10^{9}$ / 1.23 ×$\times 10^{9}$ 1.51 ×$\times 10^{1}0$ / 5.57 ×$\times 10^{9}$
Furthermore, we compared the runtime per training iteration for both ResNet and MFC-ResNet under SGD and Multiscale-SGD training strategies. Let us consider the case of $4$ levels within a Multiscale-SGD training strategy. In this case, a multiscale computation of the gradient processes for every single image on the finest resolution, 2 images on the next resolution, $4$ on the third resolution, and $8$ on the coarsest resolution. Thus, we compare its runtimes with the time it takes to process a batch of $15$ fine-scale images with the standard approach of SGD. The results are presented in Table 9. These experiments were run on an NVIDIA A6000 GPU. We observe a significant time reduction with the multiscale training strategy. Note that the MFC-ResNet is more expensive than the standard ResNet, due to the computation of the Fourier transform, which is more expensive than a simple convolution operation.
Table 9: Comparison of runtimes (milliseconds) per training iteration using standard convolution kernels and our MFCs. The results are presented for ResNet and MFC-ResNet networks with an input image size of $256\times 256$. All runs were performed on an NVIDIA A6000 GPU.
Time (ms)
Training strategy $\rightarrow$
Network $\downarrow$
SGD
Multiscale-SGD
(Ours, 4 levels)
ResNet 21.5 1.43
MFC-ResNet 27.3 1.79
@@ Appendix E Experimental setting
For the denoising task, the experiments were conducted on STL10 and CelebA datasets using networks, UNet (Ronneberger et al., 2015) and ResNet (He et al., 2016) with standard convolutions, and MFC-UNet and MFC-ResNet with mesh-free convolutions. For deblurring and inpainting tasks, the experiments were conducted on STL10 and CelebA datasets, respectively, using the same four networks as the denoising task. For the deblurring experiments, we used a blurring kernel $K(x,y)=\frac{1}{2\pi\sigma_{x}\sigma_{y}}\exp\Big{(}-\frac{x^{2}}{\sigma_{x}^{2}}-\frac{y^{2}}{\sigma_{y}^{2}}\Big{)}$ with $\sigma_{x}=\sigma_{y}=3$ to blur the input images. For the inpainting task, we introduced up to 3 (randomly chosen for each image) small rectangles with heights and widths sampled uniformly from the range $[s/12,s/6]$, where $s\times s$ is the dimension of the image. The key details of the training experimental setup for the denoising, deblurring, and inpainting tasks are summarized in Table 10. For the super-resolution task, the experiments were conducted on Urban100 dataset using networks, ESPCN (Shi et al., 2016), ResNet (He et al., 2016), MFC-ESPCN, and MFC-ResNet. The key details of the training experimental setup for the super-resolution task are summarized in Table 11.
All our experiments were conducted on an NVIDIA A6000 GPU with 48GB of memory. Upon acceptance, we will release our source code, implemented in PyTorch (Paszke et al., 2017).
Table 10: Experimental details for training for denoising, deblurring and inpainting tasks
@@
UNet: 2,537,187
ResNet: 597,699
MFC-UNet: 2,307,411
@@
SGD and Multiscale-SGD (all tasks): [2000, 2000, 2000, 2000],
Full-Multiscale-SGD (denoising/deblurring): [2000, 1000, 500, 250]
Full-Multiscale-SGD (inpainting): [2000, 2000, 2000, 2000]
Evaluation metrics MSE for denoising and deblurring, SSIM for inpainting
Table 11: Experimental details for training for super-resolution task
@@
Dataset
Urban100, consisting of paired low-resolution and high-resolution
image patches extracted for training and validation.
Network architectures
ESPCN: 5-layer super-resolution network with
64, 64, 32, 32, and 3 filters
ResNet: 9-layer ResNet-like model with 100 hidden channels per layer
MFC-ESPCN: 5-layer super-resolution network with MFCs with 32
64, 64, 32, 32, and 3 filters
MFC-ResNet: 9-layer residual network with MFCs
with 100 hidden channels per layers
Number of training parameters
ESPCN: 69,603
ResNet: 23,315
MFC-ESPCN: 80,227
MFC-ResNet: 20,755
Training strategies
SGD, Multiscale-SGD and
Full-Multiscale-SGD for all networks
Loss function Negative SSIM loss
Optimizer Adam (Kingma & Ba, 2014)
Learning rate $1\times 10^{-3}$ (constant)
Batch size strategy
Dynamic batch sizing is used, adjusting the batch size
upwards during different stages of training for
improved efficiency. For details, see Appendix C.
Multiscale levels 4
Iterations per level
SGD, Multiscale-SGD, Full-Multiscale-SGD: [250, 250, 250, 250]
Evaluation metrics SSIM
Appendix F Visualizations
@@
In this section, we include visualization of the outputs obtained from different networks with standard convolutions, as well as our MFC-based variants for various tasks such as image denoising, deblurring, inpainting, and super-resolution. Visualizations for the denoising task on the STL-10 dataset are provided in Figures 6 and 7 and on the CelebA dataset are provided in Figures 8 and 9. Visualizations for the deblurring task on the STL-10 dataset are provided in Figures 10 and 11, and for the inpainting task on the CelebA dataset are provided in Figures 12 and 13. Visualizations for the super-resolution task on the Urban100 dataset are provided in Figures 14 and 15.
Appendix G Comparing feature maps between CNNs and MFCs
@@
Figure 6: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the STL10 dataset for the denoising task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 7: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the STL10 dataset for the denoising task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 8: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the CelebA dataset for the denoising task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 9: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the CelebA dataset for the denoising task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 10: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the STL10 dataset for the deblurring task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 11: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the STL10 dataset for the deblurring task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 12: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the STL10 dataset for the inpainting task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 13: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the STL10 dataset for the inpainting task. The first two columns display the original image and data (same for all rows), followed by results from UNet, ResNet, MFC-UNet, and MFC-ResNet.
Figure 14: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the Urban100 dataset for the super-resolution task. The first column displays the low-resolution data (same for all rows), followed by results from ESPCN, ResNet, MFC-ESPCN, and MFC-ResNet.
Figure 15: A comparison of different network predictions for Single Scale, Multiscale-SGD, and Full-Multiscale-SGD for an image from the Urban100 dataset for the super-resolution task. The first two columns display the low-resolution data and original image (same for all rows), followed by results from ESPCN, ResNet, MFC-ESPCN, and MFC-ResNet.